Tony Hippwaist wants to have $742,032 in his savings account in six years. Tony opened his savings account by depositing $30,000. Tony intends to make equal deposits at the end of every three months for the next six years. Tony will earn 20% interest compounded quarterly on all deposits with the bank. Calculate the amount of each equal quarterly deposit that Tony must make in order to have $742,032 in his account in six years.
Answer #1

Solution:

Desired future value = $742,032

Quarterly interest rate = 20%/4 = 5%

Nos of quarterly periods = 6*4 = 24

Future value of initial deposit = $30,000 * (1+0.05)^24 = $30,000*3.2251 = $96,753

Future value of quarterly deposit = $742,032 - $96,753 = $645,279

Quarterly deposit amount = $645,279 / Cumulative FV factor at 5% for 24 periods

= $645,279 / 44.502 = $14,500